Jose needs an MRI (magnetic resonance imaging) scan. During the exam, Jose lies in a region of a very strong 1.6-T horizontal magnetic field pointing from Jose's waist toward his head. A medical assistant moves Jose from the scanner, reducing the magnetic field from 1.6 T to 0.1 T in 1.0 s. Consequently, the B⃗ field through Jose's 0.30-m by 0.40-m chest decreases. The conductive tissue inside his body around the outer part of his chest is a loop, with the chest as the area inside this loop.

1. Estimate the induced emf around this conducting loop as the B⃗  field decreases.

2. If the resistance of his body tissue around this loop is 5.0 Ω, what is the induced current passing around his body?

3. What is the direction of the current?
Answer #1

Given,

B1 = 1.6 T ; B2 = 0.1 T ; t = 1 s ; A = 0.3 x 0.4

1)We know that the induced emf is:

emf = -d(phi)/dt = -d(BA)/dt

emf = - A dB/dt

emf = A (B1 - B2)/t

emf = 0.3 x 0.4 x (1.6 - 0.1)/1 = 0.18 V

Hence, emf = 0.18 V

b)R = 5 Ohm

I = V/R = 0.18/5 = 0.036 A

Hence, I = 0.036 A

c)counterclockwise